OpenStudy (anonymous):

can someone tell the the steps to writing and solving an equation to find the part of a whole number 18% of 90 is what number

OpenStudy (anonymous):

18% of 90 means the same as: \[\frac{18}{100} \times \frac{90}{1}\]

OpenStudy (anonymous):

so for 5% of 522 i wrote it the same way over 1

OpenStudy (anonymous):

so the answer for the first on is 1620????

OpenStudy (anonymous):

Close, you forgot to divide by 100!

OpenStudy (anonymous):

so once i get the first number ... then i div by 100

OpenStudy (anonymous):

so its 620

OpenStudy (anonymous):

Yes, \[\frac{18}{100} \times \frac{90}{1}=\frac{18 \times 90}{100 \times 1}=\frac{1620}{100}\]

OpenStudy (anonymous):

It is not 620.

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

16.2

OpenStudy (anonymous):

16.2 is correct.

OpenStudy (anonymous):

41 of 800 is 328 right

OpenStudy (anonymous):

It is indeed correct, well done!

OpenStudy (anonymous):

question why is it always over 100

OpenStudy (anonymous):

how do you write and solvean equation to find the whole 96% of what number is 24?

OpenStudy (anonymous):

The % sign means divide by 100. "per cent" literally means "per hundred". \[\frac{96}{100} \times x = 24\]You should be able to solve for x with algebra from there.

OpenStudy (anonymous):

so 24 would b 2400???

OpenStudy (anonymous):

I think I see what you're getting at. Do you mean \[96 \times x = 2400\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then that is correct so far. What's the next step?

OpenStudy (anonymous):

x 2400 by 96

OpenStudy (anonymous):

You're close, but not quite right. If you have 96x, what do you have to do get 1x?

OpenStudy (anonymous):

96 over 1?

OpenStudy (anonymous):

96 over 1 equals 96. "over" is the right way to think though.

OpenStudy (anonymous):

so 96x 2400

OpenStudy (anonymous):

I'm afraid not.

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

What you did there was multiply by 2400.\[96 \times x = 2400\] is an equation, so whatever you do to one side, you have to do to the other side as well. If you multiply 2400 by 96, you also have to multiply 96x by 96 as well, getting you \[96 \times 96 \times x = 2400 \times 96\] which doesn't help at all!

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

You're starting with \[96 \times x = 2400\]and you want to end up with \[x=?\]To do that, you have perform some operation on both sides of the equation. To figure out what that operation is, you have to look at \[96 \times x\]and\[x\]What's different about those two numbers?

OpenStudy (anonymous):

so do you cancal out out the x and just make it 96 because u cant multi letter so u make the letter a number

OpenStudy (anonymous):

If you cancel out the x on one side, you have to do it on the other side as well. "cancelling out" in this case means dividing by x, so you go from \[96 \times x = 2400\] to \[96 = \frac{2400}{x}\] You're getting closer to the right answer, but you're not there yet!

OpenStudy (anonymous):

i think i did it wrong on my paper b/c i div 2400 by 0.96

OpenStudy (anonymous):

That would also be incorrect, yep :)

OpenStudy (anonymous):

shoot

OpenStudy (anonymous):

What's the difference between \(96x\) and \(x\)?

OpenStudy (anonymous):

96 is just plain 96

OpenStudy (anonymous):

and x is x

OpenStudy (anonymous):

25

OpenStudy (anonymous):

That's \(96x\), not 96 :)

OpenStudy (anonymous):

2400/96

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

x=25

OpenStudy (anonymous):

40% of 30 = 0.0133333333333333

OpenStudy (anonymous):

Incorrect I'm afraid. Could you post your workings?

OpenStudy (anonymous):

shoooooooot

OpenStudy (anonymous):

40/ 3000

OpenStudy (anonymous):

40/30 = 1.33333333

OpenStudy (anonymous):

\[\frac{40}{100}\times \frac{30}{1}=\frac{40 \times 30}{100 \times 1}\]

OpenStudy (anonymous):

1200 over 100

OpenStudy (anonymous):

= 12

OpenStudy (anonymous):

Yeps

OpenStudy (anonymous):

then why when you find the part it says 39% of 377 is 147.03 0.39x 377

OpenStudy (anonymous):

\[39\%=\frac{39}{100}=0.39\]They are all the same number, written differently.

OpenStudy (anonymous):

what is the difference btw finding the part of a whole and just findind the whole

OpenStudy (anonymous):

I'm only guessing here, but a part of a whole sounds like some fraction / percentage of a number.

OpenStudy (anonymous):

thanx

OpenStudy (phi):

If you need more info on this, watch Khan's video http://www.khanacademy.org/video/describing-the-meaning-of-percent?playlist=Developmental+Math He has a few on percent that are pretty good.

OpenStudy (anonymous):

wow thanx